OpenStudy (anonymous):

A 28% increase of 150.. whats the new amount?

jimthompson5910 (jim_thompson5910):

28% increase of 150 = 150 + 28% of 150 28% increase of 150 = 150 + 0.28 * 150 28% increase of 150 = ???

OpenStudy (anonymous):

192